Suppose that the Supply and Demand of maple syrup are given by Q = 100-P Q = 17 + 8 P If the government sets a binding price now p
Answer #1

Answer

Answer
the inverse demand curve is
Q=100-P
P=100-Q
the equilibrium after binding price floor is the quantity demanded at the price
P=80.39
Q=100-80.39
=19.61
Consumer surplus is the area above price and below the demand curve
CS=0.5*(Y-axis intercept of the demand curve -P)*Q
=0.5*(100-80.39)*19.61
=192.27605
=192.28
the consumer surplus is $192.28

=========

(0, 100) Price 1000 cs (19.61, 80.39) price floor (0, 80.39) 800 demand (90.778, 9.222) supply -20 20 40 60 80 quantity